OpenStudy (anonymous):

Write line equation that is perpendicular to the given line and that passes through the given point. 4x-12y=2; (10,-1)

OpenStudy (jdoe0001):

what's the slope of 4x-12y=2?

OpenStudy (anonymous):

How would I go about figuring that?

OpenStudy (jdoe0001):

solve for "y"

OpenStudy (anonymous):

1/3?

OpenStudy (jdoe0001):

yes, 1/3 sorry I was a bit away a line perpendicular to that will have a NEGATIVE RECIPROCAL slope to that what does that mean? RECIPROCAL OF 1/3 3/1 NEGATIVE OF THAT -3 then using the point-slope form \(\bf y-y_1=m(x-x_1)\) plug those values in m = slope = -3 \(\bf (x_1, y_1) = (10, -1)\)